### v2

# Dynamic Sparsity: Challenging Common Sparsity Assumptions for Learning World Models in Robotic Reinforcement Learning Benchmarks

Abstract
The use of learned dynamics models, also known as world models, can improve the sample efficiency of reinforcement learning. Recent work suggests that the underlying causal graphs of such dynamics models are sparsely connected, with each of the future state variables depending only on a small subset of the current state variables, and that learning may therefore benefit from sparsity priors. Similarly, temporal sparsity, i.e. sparsely and abruptly changing local dynamics, has also been proposed as a useful inductive bias. In this work, we critically examine these assumptions by analyzing ground-truth dynamics from a set of robotic reinforcement learning environments in the MuJoCo Playground benchmark suite, aiming to determine whether the proposed notions of state and temporal sparsity actually tend to hold in typical reinforcement learning tasks. We study (i) whether the causal graphs of environment dynamics are sparse, (ii) whether such sparsity is state-dependent, and (iii) whether local system dynamics change sparsely. Our results indicate that global sparsity is rare, but instead the tasks show local, state-dependent sparsity in their dynamics and this sparsity exhibits distinct structures, appearing in temporally localized clusters (e.g., during contact events) and affecting specific subsets of state dimensions. These findings challenge common sparsity prior assumptions in dynamics learning, emphasizing the need for grounded inductive biases that reflect the state-dependent sparsity structure of real-world dynamics.
This paper was accepted at the 40th AAAI Conference on Artificial Intelligence. This version contains the supplementary material.
1 Introduction
Reinforcement Learning (RL) promises to enable robots to learn complex, adaptive behaviors autonomously. However, it is often sample-inefficient due to the large number of environment interactions required.
Model-Based Reinforcement Learning (MBRL) addresses this issue by learning models of the environment’s dynamics, also called world models (Sutton 1988; Schmidhuber 2015; Ha and Schmidhuber 2018). World models allow agents to simulate interactions and plan effectively with fewer real-world samples (Schmidhuber 1990a, b, c, 1991; Hafner et al. 2019a, 2025). These world models learn dense dynamics where future state values are predicted based on the whole set of the current state values leading to learning spurious correlations between the states which leads to poor generalization and accuracy of prediction (Wang et al. 2021). To address this issue, recent research has proposed incorporating sparsity as an inductive bias into the learning of dynamics models for Model-Based Reinforcement Learning (MBRL), framing them as causal models (Wang et al. 2022, 2024; Lei, Schölkopf, and Posner 2024). These causal models are often assumed to be sparsely connected, meaning that each future state variable depends only on a limited subset of the current state variables and actions (Wang et al. 2022, 2024; Hwang et al. 2024; Lange and Kording 2025). Additionally, some studies advocate for modeling temporal sparsity, where dynamics change abruptly due to discrete latent transitions in the environment (Gumbsch, Butz, and Martius 2021; Jain et al. 2021; Orujlu et al. 2025).
Despite these promising directions, such assumptions are often validated only in controlled or synthetic environments, e.g. manipulation tasks with unmodifiable state variables and unmovable objects (Wang et al. 2022). It remains uncertain whether these sparsity assumptions extend from custom problems with known causal graphs to complex robotic environments featuring contacts and dynamic interactions.
In this paper, we explore the validity of these sparsity-based priors using ground-truth dynamics from the MuJoCo Playground (Zakka et al. 2025a), a suite of physics-based robotic benchmarks widely used in RL research. We focus our analysis on the sparsity of the transition functions’ Jacobians, which capture how future states linearly change with respect to current states and actions. This analysis enables the characterization of sparsity in the underlying causal structure, as the absence of causal influence corresponds to zero entries in the Jacobian. Specifically, we address four key questions:
Do the true dynamics exhibit sparsity in their causal structure?
Is such sparsity dependent on the current state?
Do the dynamics undergo temporally sparse transitions?
Does naive Multi-Layer Perceptron (MLP) training recover the ground-truth dynamics’ sparsity?
2 Related Work
This section reviews prior research most relevant to our work, beginning with general developments in continuous control environments and reinforcement learning, then narrowing our focus to methods concerned with modeling and learning environment dynamics with sparsity priors.
Continuous Control Environments:
Continuous control tasks, such as those modeled by MuJoCo (Todorov, Erez, and Tassa 2012), DeepMind Control Suite (Tassa et al. 2018), Brax (Freeman et al. 2021) and MuJoCo Playground (Zakka et al. 2025a), have become standard benchmarks for evaluating reinforcement learning (RL) algorithms due to challenging high-dimensional, continuous state and action spaces, requiring agents to learn smooth control policies for locomotion or manipulation. In this paper, we focus on the continuous control tasks shown in Figure˜1 from Mujoco Playground.
Continuous Control through RL:
Model-free approaches, such as DDPG (Lillicrap et al. 2016), PPO (Schulman et al. 2017), and SAC (Haarnoja et al. 2018), have achieved impressive results in continuous control. These methods typically rely on large amounts of interaction data, which limits their sample efficiency. In contrast, model-based RL approaches seek to improve sample efficiency by learning a predictive model of the environment’s dynamics. Notable methods include PETS (Chua et al. 2018), PlaNet (Hafner et al. 2019b), and Dreamer (Hafner et al. 2019a, 2021, 2025), which learn latent dynamics models to plan actions or improve policy learning. Efforts have also explored hybrid model-based/model-free strategies, where learned models are used to generate synthetic rollouts or augment training data. (Janner et al. 2019).
World Models and Sparse Dynamics Learning:
A growing body of work (Pitis, Creager, and Garg 2020; Wang et al. 2021, 2022, 2024; Hwang et al. 2024; Zhao et al. 2025) focuses on learning structured models of the world by combining world model learning (Ha and Schmidhuber 2018) with causal representation learning (Schölkopf et al. 2021). The main motivation for these methods is that incorporating causal structure into learning dynamics models can increase generalizability, particularly under distribution shifts or changes in controllable factors (Wang et al. 2021). Recent papers (Wang et al. 2021, 2022, 2024) propose inducing causal structure by incorporating sparsity into dynamics model learning. These papers argue that dense models, which use all current states and actions to predict future states, are prone to capturing spurious correlations between unrelated features which reduces prediction accuracy and hinders generalization. They propose learning a global context independent causal graph and assume that the local dependencies do not change over time. On the other hand, some further methods have been proposed to learn the fine-grained local context-specific independence or context (state) dependent sparsity in the causal graphs by learning local causal models (Pitis, Creager, and Garg 2020; Pitis et al. 2022; Chitnis et al. 2021; Hwang et al. 2024). These papers learn the local causal graphs for different purposes, such as data augmentation for RL agents (Pitis, Creager, and Garg 2020; Pitis et al. 2022) or exploration (Wang et al. 2023). These local causal models are learnt in various ways such as using attention scores (Pitis, Creager, and Garg 2020), examining the Jacobians of the learned dynamics model (Wang et al. 2023; Zhao et al. 2025), vector quantization of local subgraphs (Hwang et al. 2024; Zhao et al. 2025), or using hard-attention in a Transformer world model with sparsity regularisation (Lei, Schölkopf, and Posner 2024). Similarly, Orujlu et al. (2025) use RL agents to dynamically construct sparse, time-varying causal graphs, instead of the soft, dense connections typical of Transformers. Gumbsch, Butz, and Martius (2021) and Jain et al. (2021) model sparsity in the temporal evolution of latent states through the use of L0 regularization and variational sparse gating mechanisms respectively.
All the papers discussed in this section so far learn dynamics models in custom environments predominantly with known causal graphs, under the assumption that sparsity provides a useful inductive bias for MBRL in these environments. However, it was unclear if this assumption actually holds for the true dynamics of common continuous-control reinforcement learning environments as well. In this paper, we systematically investigate this assumption.
3 Background and Notation
This paper focuses on robotic RL benchmarks and in particular DM Control Suite environments (Tassa et al. 2018), as implemented in the MuJoCo Playground (Zakka et al. 2025b), for the differentiable MJX simulator a version of MuJoCo (Todorov, Erez, and Tassa 2012) written in Jax (Bradbury et al. 2018). The considered systems are vector-valued, discrete time and time-invariant, and their evolution is fully captured by where are dimensional vectors representing the ground truth states, is a dimensional vector representing the control signal (a.k.a. action), and is the simulator’s step function111Up to the indeterminism by modern GPU implementations. that calculates the next state. Figure˜2 illustrates how the states change with the actions. The observations(eg. pixel space) depend non-invertibly on the states. Performing a similar analysis on observations requires the ground-truth Jacobians between successive observations, which do not exist generally. Thus, we restrict our analysis to state space only. To interact with the environment, actions are sampled from the stochastic policy of the trained agent, i.e., . The observations capture some relevant information about the ground-truth state, but the mapping is, in general, not invertible. Note that all vectors in this manuscript are column vectors, if not stated otherwise.
Sparse Causal Graphs
The dynamics model, i.e., the function, can be viewed as a vector of scalar-valued component functions, mapping the vector valued inputs to the scalar output of each separate state dimension , i.e., . These scalar-valued output functions may depend only on a subset of the input variables where . If the output for state variable does not depend on , then no connecting edge exists in the causal graph representing the function. Sparse causal graphs (i.e., graphs with few edges) are in principle simpler to learn than dense ones, since the graph contains less information and the output can be inferred by using only a subset of the inputs. However, the knowledge of which edges are absent is generally unavailable, making it hard to harness this sparsity to simplify the learned model.
Differentiable Dynamics and Jacobians
We consider differentiable functions, and we define the first order derivatives (a.k.a. Jacobians) w.r.t. states and actions as
respectively. In other words, the Jacobians above capture the local variability of the future state given infinitesimal variations of the current state and action . The goal of this paper is to study the system’s Jacobians, which give useful information about the local interactions between variables, in order to assess the sparsity of the underlying causal graph. To this end, we collect and analyze a dataset consisting of states, actions, next states and corresponding state and action Jacobians. To make sure that we cover relevant parts of the state space, we collect the dataset by using an expert policy trained via reinforcement learning (Appendix˜A1), and using MJX, we auto-differentiate to obtain the state and action Jacobians and .
4 Using Jacobians to Assess Sparsity
To see how the Jacobians of the environment dynamics relate to sparsity in the causal graph of state variables, consider a differentiable function
where denotes the -th input variable and the component function mapping to the -th output variable . Furthermore, suppose the existence of a directed graph with vertex set
encoding direct causal relationships from inputs to outputs. An edge indicates that the output directly depends on the input . As a consequence, for each , is a function only of those inputs for which . Hence, if there is no such edge, the partial derivative of with respect to must be zero, i.e.,
Zero partial derivatives expresses that is invariant to infinitesimal changes in when is not a direct cause of in the causal graph .
Consequently, the absence of edges between certain inputs and outputs in the causal graph directly translates into sparsity in the Jacobian matrix:
More precisely, if an edge is missing in , the corresponding element of the Jacobian must be zero. In turn, the number of zero elements present in provides an upper bound on the sparsity of the causal graph. Each zero element in the Jacobian is a necessary condition for the absence of direct causal influence, but not a sufficient one, due to the possibility of higher order derivatives.
If we consider not only the derivative at a single point, but across the entire domain of we can make a stronger statement. A Jacobian element that is zero everywhere, i.e.,
implies that all higher-order derivatives with respect to are zero as well. A global zero element is thus both necessary and sufficient for the absence of a causal edge ().
5 Experiments
| Environment | Dimension |
| State | Action |
Action | |||
| BallInCup | 8 | 2 | 0 (0.00) | 0 | ||
| CartpoleBalance | 4 | 1 | 0 (0.00) | 0 | ||
| CheetahRun | 18 | 6 | 17 (5.25) | 0 | ||
| ReacherHard | 4 | 2 | 3 (18.75) | 0 | ||
| FingerSpin | 6 | 2 | 0 (0.00) | 0 | ||
| FingerTurnEasy | 6 | 2 | 0 (0.00) | 0 | ||
| WalkerRun | 18 | 6 | 17 (5.25) | 0 | ||
| SwimmerSwimmer6 | 16 | 5 | 30 (11.72) | 0 | ||
For our experimental analysis, we collect trajectories using expert reinforcement-learning policies trained with PPO (Schulman et al. 2017). To encourage exploration, we inject colored noise following the method of (Hollenstein, Martius, and Piater 2024). Additional training details are provided in Appendix˜A1. The state and action Jacobians of the ground-truth transition function are obtained via automatic differentiation. We define the sparsity value of the Jacobian matrix as the number of zeros in the Jacobian matrix divided by the total number of elements in the Jacobian matrix.
5.1 (Q1) Do the True Dynamics Exhibit Sparsity in Their Causal Structure?
This section investigates whether common Reinforcement Learning environments exhibit globally consistent sparse dynamics. We examined the Jacobians of with respect to both and for sparsity, i.e., the presence of zero elements that persist across all collected samples. As discussed in Section˜4, the condition
is both necessary and sufficient for the element of the state Jacobian to be zero.
Since we cannot exhaustively evaluate this condition, we only evaluate a necessary condition for sparsity, and provide an upper bound for the true sparsity. The results for this experiment are listed in Table˜1. We count the zero elements in both the state () and action () Jacobians. To account for floating point precision, we use a threshold of to determine whether an element is zero. Our experiments show that there are indeed environments that exhibit globally zero elements in the Jacobians for all samples tested, but that this is only the case for very few elements () and only in a handful environments (CheetahRun, ReacherHard, WalkerRun, SwimmerSwimmer6). This indicates that globally consistent sparsity is rare and is thus unlikely to be a generally important inductive bias. Instead of requiring a Jacobian element to be zero for all state and action samples, we can investigate the relaxed problem of looking at the percentage of samples where the element is zero. This is illustrated for the two environments ReacherHard and FingerTurnEasy in Figure˜3. The results clearly show the globally zero elements in the ReacherHard environment, but further highlight that specific elements can be zero only for a certain percentage of the samples, as for FingerTurnEasy. This hints at a different, potentially more widely applicable sparsity assumption: state dependent sparsity, which we investigate next.
5.2 (Q2) Is Such Sparsity Dependent on the Current State?
In the previous section, we examined globally zero elements in the Jacobians. We found that some environments do have such zero elements, but most do not. Even when they do, there are only a few of these globally zero elements. However, relaxing the requirement of global sparsity for the Jacobian elements, we can investigate partial sparsity in the Jacobian, i.e., when the Jacobian elements are zero only for fraction of the samples.
Sparsity (i.e., zero elements) in the Jacobian for a fraction of the sampled states is a necessary condition for state-dependent sparsity, which, as indicated by Figure˜3, is present in at least some of the environments. In this section, we turn to looking at state-dependent sparsity in more detail. Figure˜4 illustrates state dependent sparsity for the FingerSpin environment, through a t-SNE embedding of the state, action and next state (s, a, s’) tuple and color coding the combined state and action Jacobian’s sparsity at each point. The figure illustrates a sequence of frames, the corresponding transitions through the t-SNE embedded state-space and the sparsity values. As can be observed in the first frames, the sparsity is lower, when the finger interacts with the object. This sequence of steps with low sparsity followed by few steps of high sparsity, repeats as the object briefly comes into contact with the finger (=low sparsity) and then spins freely (=high sparsity). Sparsity could be induced in a learned representation, e.g, through regularization (similar to Lei, Schölkopf, and Posner (2024). However, this method requires tuning the amount of sparsity that is induced. Figure˜5 illustrates histograms of the observed sparsity across all samples for all tested environments. The histograms indicate that the sparsity only assumes certain values, presumably based on whether the agent touches the ground (CheetahRun, WalkerRun), or interacts with the object (BallInCup, FingerSpin, FingerTurnEasy). The figure also illustrates the mean-sparsity for each environment. Presumably, if sparsity regularization is used, it should be tuned to induce a similar amount of sparsity on average. While Figure˜5 indicates that the sparsity only assumes specific values, it is unclear how the sparsity changes temporally.
5.3 (Q3) Do the Dynamics Undergo Temporally Sparse Transitions?
The previous section established that the system assumes specific sparsity values, likely due to the structure of the interacting objects and the robot, such as objects coming into contact or airborne phases of the gait cycle. If sparsity relates to the gait cycle or contact events, its behavior is expected to be temporally consistent—remaining stable over multiple steps before switching—which can be observed in Figure˜4. The system’s dynamics can be described by differential equations where changes in contact correspond to transitions between different underlying differential equations. Therefore, incorporating inductive biases favoring modelling of temporally sparse switches between underlying dynamics may improve the learned performance (Gumbsch, Butz, and Martius 2021; Jain et al. 2021).
Gumbsch, Butz, and Martius (2021) have pursued a similar approach, employing a temporally sparsely changing recurrent network to model both agent behavior in partially observable environments and environment dynamics. In their method, a hyper-parameter penalizes the changes and thus implicitly regulates how often the system switches between states, or inversely, stays in the same state. In this section, we examine whether sparsity changes occur in a temporally sparse manner and analyze the distribution of durations for which these sparsity states persist. Figure˜6 depicts the evolution of the percentage of zero elements over time during a single rollout with the expert agent, using the stochastic Gaussian policy. In the BallInCup environment, a ball tethered to a controllable cup by a string must be caught by moving the cup. Sparsity is high when the string is loose and low when the ball is inside the cup or the string is taut, indicating temporally sparse changes. Figure˜7 presents the distribution of durations, measured by consecutive timesteps, during which Jacobian elements remain in the close to zero or non-zero state. The duration of elements which remain zeros for the maximum length of the episode are not shown. Many of the elements in these Jacobians remain in the same state for durations of multiple steps, hinting at underlying structure that sparsely changes in time. The mean duration of change likely poses as a useful default hyperparameter for methods attempting to regularize for temporally sparse switching.
5.4 (Q4) Does Naive MLP Training Recover the Ground-truth Dynamics’ Sparsity?
The previous sections demonstrated that even though global sparsity is rare but sparsity does occur in the ground-truth dynamics, though only to a limited, state-dependent extent.
A natural follow-up question is whether a multi-layer perceptron (MLP) trained to approximate can capture this sparsity. To investigate this question, we tested datasets collected from our benchmark environments. Specifically, we trained a two-layer neural network with ELU activations and 512 units per layer (Appendix˜A2), using mean squared error (MSE) loss to predict the next state. Datasets were normalized before training (Section˜A2.2). We consider a Jacobian entry of the MLP with an absolute value below a threshold of to be effectively zero.
In addition to the baseline MSE loss, we tested the separate inclusion of ground-truth Jacobian losses (MSE and mean absolute error (MAE)) and an L1 penalty applied exclusively to the predicted Jacobians. We also introduced a sparsity-aware error (SAE) loss, which applies an MAE loss only to the Jacobian elements that are expected to be zero based on the ground-truth Jacobian. The results, summarized in Figure˜8, revealed a slight trend: using an MAE loss for the Jacobian appeared to improve next-state prediction performance, while MSE loss slightly reduced it. However, these differences were not significant. In contrast, the induced sparsity increased significantly when ground-truth Jacobian losses were included, but still remained multiple orders of magnitude below the target sparsity. Notably, the SAE loss increased sparsity without degrading prediction performance, whereas applying L1 regularization to the Jacobians increased sparsity but negatively impacted state prediction accuracy. A reduced sparsity value indicates more entangled predictions compared to the ground truth, and consequently higher prediction errors from spuriously learned correlations. While the naive MLP recovers some sparsity in its predictions, it is insufficient to fully capture the ground-truth sparsity, indicating a need for improved world model architectures.
6 Discussion
Main Findings
Our study provides insights into the role of sparsity in learning the dynamics of classical reinforcement learning (RL) environments and its implications for model-based reinforcement learning (MBRL). First, we observed that globally sparse causal structures, as indicated by consistently zero Jacobian elements, are rare across most environments: While some environments, such as ReacherHard and CheetahRun, exhibited limited global sparsity, the majority showed dense interactions between state and action variables. This suggests that enforcing strong global sparsity priors in learned dynamics models may not be universally beneficial. Interestingly, we found evidence of state-dependent sparsity i.e. the causal structure of the dynamics changes based on the current state. For example, in the FingerSpin environment, sparsity was higher when the finger was not in contact with the object and lower during interactions.
In our learning experiments, we observed that while a two-layer MLP trained to approximate could recover some sparsity in its predictions, the induced sparsity remained far below the target sparsity. Using MAE-based Jacobian losses slightly improved next-state prediction performance compared to MSE-based losses, though the differences were not statistically significant. Including ground-truth Jacobian losses increased induced sparsity but did not fully capture the sparsity present in the ground-truth dynamics, leaving the network’s predictions more entangled. This likely leads to reduced generalization and more compounding errors during multistep prediction.
Limitations
While this study provides valuable insights, there are limitations which provide room for further improvement. Due to computational budget reasons the experiments were conducted with a limited set of agents and environments. Increasing the set of agents and environments would further generalize the findings. Similarly, the architectural space is vast and architectural exploration was thus limited to the most intuitive choice, i.e., equal width MLPs. While expert trajectories are arguably the most relevant data for modelling the environment while also achieving good task performance—the collected data were limited to stochastic-policy rollouts of trained agents—more diverse data could further validate the results.
7 Conclusion
While sparsity priors hold promise for improving sample efficiency and generalization, their effectiveness depends on alignment with the true structure of the environment. Overly strong or misaligned priors could hinder learning by blocking important interactions. Our findings indicate that sparsity is indeed present in many reinforcement learning environments, but requires modelling in a state dependent way. Additionally, we observed that changes in sparsity often occur in a temporally sparse manner, with periods of stable sparsity interspersed with abrupt transitions such as contact events or phase changes in locomotion. Our results also indicate that naive MLP implementations are insufficient to fully capture and exploit these sparsity structures. These findings highlight the need for further development of dynamics model architectures that can explicitly model sparsity and thus dynamically adapt their causal structure based on state and time, thereby improving generalization and interpretability.
Acknowledgements
We thank Christof Beck, Ananth Rachakonda, and Henri Geiß for their helpful comments on earlier drafts.
This research was partially funded by the Austrian Science Fund (FWF):
I 5755-N (ELSA), and by
the Autonomous Province of Bolzano-Bozen - South
Tyrol under Funding Agreement 10/2024, Abstractron.
Supplementary Appendix
Appendix A1 Agents & Data Collection
We generate a dataset of trajectories per environment using pre-trained PPO agents based on the implementation by (Raffin et al. 2021), more specifically following the colored noise methodology proposed by (Hollenstein, Martius, and Piater 2024). Multiple agents were trained using the hyperparameter configurations described in Table˜A1.2. For each environment, the best performing agent was selected. The mean reward achieved across the full episodes in our dataset is shown in Table˜A1.2.
Dataset generation then uses the selected best agents. Dataset generation is necessary because there are no publicly available datasets containing ground-truth Jacobians of the transition functions. For the dataset, we collect a total of million steps per environment, using the learned stochastic agent policy (i.e., sampling actions from the parameterized-Gaussian policy). To improve fidelty MJX and JAX where switched to float64 mode. The MLP parameters in the learning experiments were represented in float32.
| Param. | Value | Default |
| lr | 5e-05 | 0.0003 |
| n-steps | 64 | 2048 |
| batch-size | 64 | |
| n-epochs | 10 | |
| gamma | 0.99 | |
| gae-lambda | [0.0, 0.95] | 0.95 |
| clip-range | 0.2 | |
| normalize-advantage | True | |
| ent-coef | 0 | |
| vf-coef | 0.5 | |
| use-sde | False | |
| sde-sample-freq | -1 | |
| max-grad-norm | 0.5 | |
| stats-window-size | 100 | |
| eval-episodes | 50 | |
| eval-freq | 10240 | |
| n-envs | 128 | |
| total-timesteps | 2.04801e+06 | |
| seed | [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] | |
| noise-color | [0.0, 0.5, 1.0, 2.0] | |
| n-eval-envs | 10 |
| Environment | Mean Reward(Evaluation) |
| CartpoleBalance | 999.88 |
| BallInCup | 960.00 |
| FingerSpin | 902.56 |
| ReacherHard | 709.20 |
| SwimmerSwimmer6 | 650.36 |
| FingerTurnEasy | 515.66 |
| CheetahRun | 351.75 |
| WalkerRun | 327.80 |
Appendix A2 Learning Experiments
| Experiment Parameters | |
| activation-function | elu |
| batch-size | 512 |
| dropout-rate | 0.1 |
| env | BallInCup, CartpoleBalance, CheetahRun, FingerSpin, FingerTurnEasy, ReacherHard, SwimmerSwimmer6, WalkerRun |
| epochs | 100 |
| initializer | kaiming_normal |
| layer-depth | 2 |
| layer-width | 512 |
| learning-rate | 0.002 |
| loss-mode | state, state-jacobian, state-jacobian-l1, state-jacobian-l1-regularizer, state-jacobian-l1-sparse-only |
| normalize-data | True |
| randomize-steps | True |
| seed | 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15 |
| test-split | 0.1 |
| weight-decay | 0.001 |
A2.1 Hyperparameter Search for Dynamics Models
All hyperparameter search runs use a batch size of , are conducted across seeds, and otherwise use the same hyperparameters as shown in Table˜A2.1, unless specified otherwise.
Search Run: 1
We run a range of experiments across the hyperparemeter selections outlined in Table˜A2.2. All runs are ranked per environment according to the final state prediction loss on the test set. For determining our final choice per hyperparmameter, we determine the mean rank across environments and marginalize across all other hyperparameters. The resulting ranks and choices are shown in Tables˜A2.3, A2.4, A2.5 and A2.6.
| Hyperparameter | Values |
| learning-rate | [0.01, 0.0001, 0.001] |
| layer-width | [512, 1024] |
| layer-depth | [4, 6, 2] |
| dropout-rate | [0.1, 0.5, 0.3] |
| learning-rate | Test Rank |
| 0.001 | 338.43 |
| 0.0001 | 362.32 |
| 0.01 | 515.75 |
| layer-width | Test Rank |
| 512 | 396.23 |
| 128 | 399.47 |
| 32 | 420.80 |
| layer-depth | Test Rank |
| 2 | 245.32 |
| 4 | 437.56 |
| 6 | 533.62 |
| dropout-rate | Test Rank |
| 0.10 | 237.48 |
| 0.30 | 415.37 |
| 0.50 | 563.64 |
Search Run: 2
We repeat this procedure with the hyperparameter ranges shown in Table˜A2.8. All remaining hyperparameters are respectively fixed to the initial values, or to the values found in the first search. The resulting ranks and choices are shown in Tables˜A2.9, A2.10 and A2.11.
| Hyperparameter | Values |
| layer-width | [512, 1024] |
| layer-depth | [3, 2] |
| weight-decay | [0.01, 0.001] |
| layer-width | Test Rank |
| 512 | 19.75 |
| 1024 | 21.25 |
| layer-depth | Test Rank |
| 2 | 16.61 |
| 3 | 24.39 |
| weight-decay | Test Rank |
| 0.00 | 20.38 |
| 0.01 | 20.62 |
Search Run: 3
We perform the same procedure a final time to find a choice of activation function between the Exponential Linear Units (ELU) (Clevert, Unterthiner, and Hochreiter 2016) and the Rectified Linear Unit (ReLU) (Fukushima 1969). The resulting ranks and choice are shown in Table˜A2.13.
| activation-function | Test Rank |
| elu | 8.05 |
| leaky-relu | 12.80 |
For the learning experiments we double both batch size and learning rate to better utilize our hardware. We end up with the final hyperparameter choices specified in Table˜A2.1.
A2.2 Data Normalization
We apply Z-score normalization to all variables (states, actions, Jacabions, etc.). For a vector :
The unnormalized value is recovered as:
For a mapping (e.g., from actions to next-states), define corresponding mean and standard deviations , . The normalized mapping is:
Let (unnormalized Jacobian), and (normalized Jacobian). Then:
or elementwise,
with scaling matrix , i.e.,
A2.3 Prediction Loss per Environment
A2.4 Sparsity per Environment
Appendix A3 Sparsity Plots
A3.1 Sparsity Heatmaps Showing Zero-Duration Proportions
Sparsity heatmaps showing the duration for which the elements in the state and action Jacobians were zero as a proportion of the total environment steps in a rollout. The following heatmaps are for environments which were not included in the paper for both state and action Jacobians in the Figure˜A3.1 and Figure˜A3.2
A3.2 Sparsity time evolution plots
The time evolution of sparsity of state and action Jacobians for the environments not included in the paper are provided here in Figure˜A3.3.
A3.3 Code and Dataset for training agents
The code is part of files submitted as part of the supplementary. The code includes the training of the MLPs and the data collection scripts. (partial code released).
Example datasets and their descriptions can be found at this link - https://doi.org/10.5281/zenodo.16743573